# Exploring Adapter Design Tradeoffs for Low Resource Music Generation

### v1

@@ Abstract.
Fine-tuning large-scale music generation models, such as MusicGen and Mustango, is a computationally expensive process, often requiring updates to billions of parameters and, therefore, significant hardware resources. Parameter-Efficient Fine-Tuning (PEFT) techniques, particularly adapter-based methods, have emerged as a promising alternative, enabling adaptation with minimal trainable parameters while preserving model performance. However, the design choices for adapters, including their architecture, placement, and size, are numerous, and it is unclear which of these combinations would produce optimal adapters and why, for a given case of low-resource music genre. In this paper, we attempt to answer this question by studying various adapter configurations for two AI music models, MusicGen and Mustango, on two genres: Hindustani Classical and Turkish Makam music.
Our findings reveal distinct trade-offs: convolution-based adapters excel in capturing fine-grained local musical details such as ornamentations and short melodic phrases, while transformer-based adapters better preserve long-range dependencies crucial for structured improvisation. Additionally, we analyze computational resource requirements across different adapter scales, demonstrating how mid-sized adapters (40M parameters) achieve an optimal balance between expressivity and quality. Furthermore, we find that Mustango, a diffusion-based model, generates more diverse outputs with better adherence to the description in the input prompt while lacking in providing stability in notes, rhythm alignment, and aesthetics. Also, it is computationally intensive and requires significantly more time to train. In contrast, autoregressive models like MusicGen offer faster training and are more efficient, and can produce better quality output in comparison, but have slightly higher redundancy in their generations.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Applied computing Sound and music computingccs: Information systems Multimedia content creationccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Figure 1. Adapter-based fine-tuning: Exploring different placements and architecture types (fully connected linear layers, Convolution, Transformer) in an encoder-decoder architecture.
The contributions of our paper are threefold [^1] We plan to release our code and trained models after the paper is accepted to encourage further research in this direction.:
@@
    • Architecture: We systematically compare fully connected (linear) layer-based, Convolution-based, and Transformer-based adapters, analyzing their impact on generation quality and computational efficiency.
@@ 2. Parameter-Efficient Fine-Tuning with Adapters
In this section, we discuss the design considerations for integrating adapters into music generation models. While adapter-based Parameter-Efficient Fine-Tuning (PEFT) techniques offer significant computational benefits, they also introduce trade-offs that need to be balanced, primarily between quality and parameter efficiency. Specifically, the challenges in music generation, such as capturing long-range dependencies, multi-scale structures, and genre-specific feature,s add complexity to this optimization.
@@ 2.1. Model Selection
We employ MusicGen (Copet et al., 2023) and Mustango (Melechovsky et al., 2024) to explore cross-genre adaptation in music generation. MusicGen is a Transformer-based autoregressive model that can generate music from text prompts and from partial melodies being prompted using EnCodec representations, excelling in style transfer and offering high controllability. In contrast, Mustango extends pre-trained language models for text-to-music synthesis by leveraging a diffusion model, which conditions on text prompts, chord progressions, and beat information. This distinction allows us to analyze how different architectures adapt to underrepresented non-Western genres, providing valuable insights into adapter placement and training methodologies for distinct model types.
@@ 2.2. Trade-offs between Efficiency and Quality
  • Efficiency: Adapters drastically reduce the number of parameters that need to be fine-tuned, allowing for faster training and lower computational requirements. In our experiments, we explore adapter sizes ranging from 2M to 70M parameters. By limiting the fine-tuning to a small portion of the model, adapters help make large-scale models more accessible for tasks requiring fewer computational resources.
@@ 2.3. Adapter Architecture Design Choices
Music generation poses unique challenges that influence how adapters should be designed and placed within the model. For Hindustani Classical (Jairazbhoy, 1971) and Maqam (Signell, 2008) music, adapter-based fine-tuning must address unique challenges such as complex melodic progressions, microtonal nuances, and long-form structures. Fully connected (linear) layer-based adapters offer efficiency but may struggle to capture the local temporal or hierarchical dependencies inherent in specific musical genres or traditions, making CNNs a better choice for these tasks. Convolution (CNN)-based adapters can model short-term dependencies effectively, capturing local patterns in music. However, they struggle to represent extended melodic developments, which require long-range dependencies that are vital in many musical traditions. In contrast, Transformer-based adapters are designed to capture long-term dependencies by leveraging self-attention mechanisms that allow them to model relationships across the entire sequence. This makes Transformers more effective than CNNs for capturing extended melodic phrasing and intricate, evolving musical structures, which are crucial in complex compositions. However, they may struggle with modeling short-term or local dependencies, which can impact the quality of finer, localized musical details, while also being more computationally resource-intensive during training. We experiment with three different adapter architectures, with Figure 1 illustrating the key components of each adapter type.
@@
  • Fully connected (linear) layer-based: Linear layers (Houlsby et al., 2019) serve as efficient feature extractors, enabling compact representations of musical sequences. However, they struggle with long-range dependencies, which are crucial for raga and maqam progression. In our setup, this adapter compresses the input sequence into a low-dimensional bottleneck space using a down-projection layer, applies GELU (Hendrycks and Gimpel, 2016) activation, and restores the original sequence length through an up-projection layer. The choice of the bottleneck dimension plays a key role in determining the trainable parameter count, as a larger bottleneck would increase the number of parameters in both projection layers. A dropout layer prevents overfitting, and residual connections ensure the retention of key musical features. For MusicGen, to maintain compatibility with the stereo (2-channel) format, the output is expanded accordingly. For Mustango, this adapter type is not used, as it follows a UNet architecture with inputs having more than three dimensions, making only CNN and Transformer architectures suitable for use.
@@
  • Convolution-Based (CNN): Convolution neural network (LeCun and Bengio, 1998) adapters are effective in capturing local dependencies and fine-grained features such as gamakas, meends, and murkis key ornamentations in Hindustani Classical and Maqam music. In our setup, this adapter begins with a down-projection convolutional layer that reduces the input dimensionality while preserving sequence length. A deep residual bottleneck module follows, consisting of multiple dilated residual blocks (Yu et al., 2017) that enable the model to process both short-term and long-term dependencies in the music signal. The bottleneck dimension directly influences the number of trainable parameters, particularly in convolutional layers and residual blocks. Additionally, a Squeeze-and-Excitation (SE) block (Hu et al., 2018) applies channel-wise attention, allowing the adapter to dynamically focus on salient musical features. The up-projection layer restores the original sequence dimensions, ensuring minimal disruption to the pre-trained model. A dropout layer is incorporated to prevent overfitting.
@@ 2.4. Placement Strategies for Adapters
In our placement experiments, shown in Figure 1, we tested inserting adapters into both the middle and late layers of MusicGen and Mustango. For each configuration, we listened to the generated audio and evaluated it using standard metrics. We trained these models for more than 10 epochs, listening to the quality of the generated audio to assess the output produced by the model.
@@ 2.4.1. MusicGen
For MusicGen, we found that placing adapters in the middle layers often led to a complete breakdown in generation. The audio was not just of low quality; it was severely distorted, often consisting of loud beeping sounds or static, with no recognizable musical structure. This failure occurred even for prompts where the original base models could generate coherent, high-quality music.
@@ 2.4.2. Mustango
For Mustango, we incorporated adapters into the UNet architecture, which consists of three distinct blocks: the downsampling block, the middle block, and the upsampling block. Each of these blocks contains a combination of ResNet and Transformer layers, repeated multiple times. We experimented with different adapter placements within these blocks, testing configurations where adapters were added after each Transformer layer, after each ResNet layer, and after each complete block.
Our initial experiments showed that adding adapters after the Transformer layers resulted in output that lacked structure and essentially sounded like random noise. Through further analysis, we discovered that the most effective configuration was to place adapters after each complete block. This arrangement allowed for more coherent output, maintaining the model’s structure while improving its adaptation to the target genre.
Interestingly, this pattern was consistent across both models and genres. Whether for Hindustani Classical or Turkish Makam, inserting adapters in the middle layers caused similar degradation. Our hypothesis is that the middle layers are critical for encoding the model’s core understanding of music, such as timbre, harmonic progression, and rhythm. Modifying these internal activations appears to destabilize the learned flow of information, causing the model to “forget” or corrupt its prior musical understanding. These findings suggest that respecting the internal structure of generative music models is critical during fine-tuning. Intervening too early in the computation pipeline can disrupt key learned representations. In contrast, late-layer adapters allow for specialization and stylistic adaptation without interfering with the foundational knowledge encoded in earlier layers.
@@ 3.1.2. Prompt Formation
(a) Subjective quality & aesthetic evaluation for Hindustani Classical
(b) Subjective quality & aesthetic evaluation for Turkish Makam
Figure 3. Human evaluation for Hindustani Classical & Turkish Makam musical genres.
For training text-to-music models, we have to generate text prompts for each equivalent audio sample. To make sure each prompt is descriptive, we utilise the above metadata combinations for each audio sample and fit them into five pre-defined prompt templates. This enables us to give separate descriptive entities to the model, ensuring that the model does not overfit while giving genre-specific entities to recognise and generate music accordingly.
@@ 3.3.2. Subjective Evaluation
For subjective human evaluation, we used a similar framework as described in  Mehta et al. (2025) by implementing an arena-style setup and an elimination-based strategy to compare the generative performance of our models subjectively. In this approach, for each of the 10 distinct prompts, we generated audio snippets from two models at a time and compared them on the basis of clarity, auditory pleasantness, and freedom from unwanted artifacts. The model that wins on a majority of these prompt-wise comparisons advances to the next round, and this process is repeated until a single model remains. We compare five models (three from MusicGen and two from Mustango) for both Hindustani and Makam, resulting in 40 total matchups per genre. Ultimately, the best-performing model is identified as the winner for that genre.
@@ 4.1. Objective Evaluations
Figure 2 (left plots) shows results for MusicGen for both Hindustani Classical and Turkish Makam music. Across all configurations, we find that the 40M parameter scale offers the best trade-off between musical quality and computational cost. At this scale, FAD and FD scores stabilize at low values across architectures and genres, while training time remains significantly lower than that of the larger 70M models. For example, Hindustani_MG-C and Makam_MG-T at 40M both achieve excellent FAD scores (10.0 and 13.0, respectively), indicating good audio quality, without incurring the diminishing returns observed beyond this scale. This sweet spot can be attributed to the nature of adapter-based PEFT; small adapters (e.g., 2M or 8M) lack sufficient capacity to capture complex musical dependencies, conversely, very large adapters (e.g., 70M) not only increase memory and compute requirements but also risk overfitting or disrupting the model’s learned representations particularly noticeable in MG-L configurations, where FD scores for both genres degrade sharply at 70M. Among architectures, convolution-based adapters (MG-C) performs best for Hindustani classical, while transformer-based adapters (MG-T) perform best for Turkish Makam.
Figure 2 (right plots) shows Mustango results for both Hindustani Classical and Turkish Makam music. Unlike MusicGen, whose best trade-off emerged at the 40M scale Mustango demonstrates its most stable performance in the 20M–40M parameter range. At smaller adapter sizes (2M or 8M), model capacity is insufficient for capturing complex musical structures, reflected in higher FAD scores (e.g., Hindustani_MS-C at 2M hovers near 12). In contrast, the largest 70M adapters yield inconsistent improvements and can even degrade fidelity metrics (e.g., FD for Makam_MS-T climbs above 100 at 70M), suggesting a risk of overfitting or destabilizing previously learned representations. In particular, 40M adapters typically achieve low FAD values, especially for Hindustani CNN (MS-C) at 6.4 and Makam CNN (MS-C) at 8.39. Among architectures, convolution-based adapters (MS-C) performs best for both Hindustani classical and Turkish Makam in the case of .
Both MusicGen and Mustango share the common trend that mid-range adapter sizes avoid the trade-offs of too-little capacity (leading to higher FAD/FD scores) and too-large capacity (increased risk of overfitting and heavier compute demands). In MusicGen, the optimal point holds firmly at 40M; in Mustango, peak performance is often achieved slightly earlier, around 20M parameters, though 40M remains competitive. In both models, architecture preferences are genre-dependent, with CNN adapters often favored for Hindustani classical and Transformer adapters for Turkish Makam.

### v2

@@ Abstract.
Fine-tuning large-scale music audio generation models, such as MusicGen and Mustango, is a computationally expensive process, often requiring updates to billions of parameters and, therefore, significant hardware resources. Parameter-Efficient Fine-Tuning (PEFT) techniques, particularly adapter-based methods, have emerged as a promising alternative, enabling adaptation with minimal trainable parameters while preserving model performance. However, the design choices for adapters, including their architecture, placement, and size, are numerous, and it is unclear which of these combinations would produce optimal adapters and why, for a given case of low-resource music genre. In this paper, we attempt to answer this question by studying various adapter configurations for two AI music models, MusicGen and Mustango, on two genres: Hindustani Classical and Turkish Makam music.
Our findings reveal distinct trade-offs: convolution-based adapters excel in capturing fine-grained local musical details such as ornamentations and short melodic phrases, while transformer-based adapters better preserve long-range dependencies crucial for structured improvisation. Additionally, we analyze computational resource requirements across different adapter scales, demonstrating how mid-sized adapters (40M parameters) achieve an optimal balance between expressivity and quality. Furthermore, we find that Mustango, a diffusion-based model, generates more diverse outputs with better adherence to the description in the input prompt while lacking in providing stability in notes, rhythm alignment, and aesthetics. Also, it is computationally intensive and requires significantly more time to train. In contrast, autoregressive models like MusicGen offer faster training and are more efficient, and can produce better quality output in comparison, but have slightly higher redundancy in their generations. We release our datasets, models and training code in the following github repository: Github.
@@
journalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755766isbn: 979-8-4007-2035-2/2025/10ccs: Applied computing Sound and music computingccs: Information systems Multimedia content creationccs: Computing methodologies Artificial intelligence
@@ 1. Introduction
Figure 1. Adapter-based fine-tuning: Exploring different placements and architecture types (fully connected layers, Convolution, Transformer) in an encoder-decoder architecture.
The contributions of our paper are threefold:
@@
    • Architecture: We systematically compare fully connected layer-based, Convolution-based, and Transformer-based adapters, analyzing their impact on generation quality and computational efficiency.
@@ 2. Parameter-Efficient Fine-Tuning with Adapters
In this section, we discuss the design considerations for integrating adapters into music generation models. While adapter-based Parameter-Efficient Fine-Tuning (PEFT) techniques offer significant computational benefits, they also introduce trade-offs that need to be balanced, primarily between quality and parameter efficiency. Specifically, the challenges in music generation, such as capturing long-range dependencies, multi-scale structures, and genre-specific features add complexity to this optimization.
@@ 2.1. Model Selection
We employ MusicGen (Copet et al., 2023) and Mustango (Melechovsky et al., 2024) to explore cross-genre adaptation in music audio generation. MusicGen is a Transformer-based autoregressive model that can generate music from text prompts and from partial melodies being prompted using EnCodec representations, excelling in style transfer and offering high controllability. In contrast, Mustango extends pre-trained language models for text-to-music synthesis by leveraging a diffusion model, which conditions on text prompts, chord progressions, and beat information. This distinction allows us to analyze how different architectures adapt to underrepresented non-Western genres, providing valuable insights into adapter placement and training methodologies for distinct model types.
@@ 2.2. Trade-offs between Efficiency and Quality
  • Efficiency: Adapters drastically reduce the number of parameters that need to be fine-tuned, allowing for faster training and lower computational requirements. In our experiments, we explore adapter sizes ranging from 2M to 70M parameters. By limiting the fine-tuning to a small portion of the model, ranging from 0.1% to 5% of the base model, adapters help make large-scale models more accessible for tasks requiring fewer computational resources.
@@ 2.3. Adapter Architecture Design Choices
Music generation poses unique challenges that influence how adapters should be designed and placed within the model. For Hindustani Classical (Jairazbhoy, 1971) and Maqam (Signell, 2008) music, adapter-based fine-tuning must address unique challenges such as complex melodic progressions, microtonal nuances, and long-form structures. Fully connected layer-based adapters offer efficiency but may struggle to capture the local temporal or hierarchical dependencies inherent in specific musical genres or traditions, making CNNs a better choice for these tasks. Convolution (CNN)-based adapters can model short-term dependencies effectively, capturing local patterns in music. However, they struggle to represent extended melodic developments, which require long-range dependencies that are vital in many musical traditions. In contrast, Transformer-based adapters are designed to capture long-term dependencies by leveraging self-attention mechanisms that allow them to model relationships across the entire sequence. This makes Transformers more effective than CNNs for capturing extended melodic phrasing and intricate, evolving musical structures, which are crucial in complex compositions. However, they are more computationally resource-intensive during training and require more data for training and adaptation. We experiment with three different adapter architectures, with Figure 1 illustrating the key components of each adapter type.
@@
  • Fully connected layer-based: Fully connected layers (Houlsby et al., 2019) serve as efficient feature extractors, enabling compact representations of musical sequences. However, they are not as good as transformers and cnn architectures in capturing long-range dependencies and time sequence modelling which are crucial for raga and maqam progression. In our setup, this adapter compresses the input sequence into a low-dimensional bottleneck space using a down-projection layer, applies GELU (Hendrycks and Gimpel, 2016) activation, and restores the original sequence length through an up-projection layer. The choice of the bottleneck dimension plays a key role in determining the trainable parameter count, as a larger bottleneck would increase the number of parameters in both projection layers. A dropout layer prevents overfitting, and residual connections ensure the retention of key musical features. For MusicGen, to maintain compatibility with the stereo (2-channel) format, the output is expanded accordingly. For Mustango, this adapter type is not used, as it follows a UNet architecture with inputs having more than three dimensions, making only CNN and Transformer architectures suitable for use.
@@
  • Convolution-Based (CNN): Convolution neural network (LeCun and Bengio, 1998) adapters are effective in capturing local dependencies and fine-grained features which can include gamakas [^1] Gamaka can be understood as embellishment done on a note or between two notes., meends [^2] In Hindustani music, meend refers to a glide from one note to another., and murkis [^3] Murki is a short taan or inverted mordent in Hindustani classical music. key ornamentations in Hindustani Classical and Maqam music. In our setup for MusicGen, the adapter begins with a down-projection convolutional layer that reduces the input dimensionality while preserving sequence length. A deep residual bottleneck module follows, consisting of residual blocks (Yu et al., 2017) that enable the model to process both short-term and long-term dependencies in the music signal. The bottleneck dimension directly influences the number of trainable parameters, particularly in convolutional layers and residual blocks. Additionally, a Squeeze-and-Excitation (SE) block is added in MusicGen which (Hu et al., 2018) applies channel-wise weight, allowing the adapter to dynamically focus on salient musical features. The up-projection layer restores the original sequence dimensions, ensuring minimal disruption to the pre-trained model. A dropout layer is incorporated to prevent overfitting.
@@ 2.4. Placement Strategies for Adapters
In our placement experiments, shown in Figure 1, we tested inserting adapters into both the middle and late layers of MusicGen. For Mustango, we performed similar experiments on intermediate vs final layer placement and UNet block-wise addition of adapters. For each configuration, we listened to the generated audio and evaluated it using standard metrics. We trained these models for more than 10 epochs, listening to the quality of the generated audio to assess the output produced by the model.
@@ 2.4.1. MusicGen
For MusicGen, we found that placing adapters in the middle layers often led to a complete breakdown in generation. The audio was not just of low quality; it was severely distorted, often consisting of loud beeping sounds or static, with no recognizable musical structure. This failure occurred even for prompts where the original base models could generate coherent, high-quality music. We hypothesize that such issues arise due to lack of data to finetune these layers in comparison to the data used for pre-training.
@@ 2.4.2. Mustango
For Mustango, we incorporated adapters into the UNet architecture used for denoising, which consists of three main blocks: downsampling, mid-sampling, and upsampling. Each block contains multiple ResNet and Transformer layers. We experimented with various adapter placements after each transformer layer, after each ResNet layer, and after each complete block.
Our initial experiments showed that inserting adapters after individual Transformer layers resulted in outputs that lacked structure and often resembled random noise. Further analysis revealed that placing adapters after each complete block was far more effective. This configuration preserved the internal flow of information while allowing the model to adapt meaningfully to the target genre.
This pattern held consistently across both MusicGen and Mustango, and for both Hindustani Classical and Turkish Makam music. Inserting adapters into intermediate layers—such as within the encoder/decoder of MusicGen or within the internal layers of Mustango—led to degraded outputs. We hypothesize that these layers encode foundational musical concepts like timbre, harmonic progression, and rhythm. Modifying these activations disrupts the model’s core understanding, causing it to forget or corrupt previously learned representations.
Finally, we observed that removing adapters from any block distorted learning, suggesting that adapter presence at the end of each block—downsampling, mid-sampling, and upsampling—is crucial for stability. These findings underscore the importance of adapter placement and highlight the value of late-layer interventions in preserving structure while enabling stylistic adaptation.
@@ 3.1.2. Prompt Formation
Figure 3. Human evaluation of subjective quality and aesthetics for (a) Hindustani Classical and (b) Turkish Makam music.
For training text-to-music models, we have to generate text prompts for each equivalent audio sample. To make sure each prompt is descriptive, we utilise the above metadata combinations for each audio sample and fit them into five pre-defined prompt templates. Each prompt is a paraphrased version of the same base, with blanks for melodic line, instruments, and rhythmic pattern. This augmentation promotes diversity and helps prevent overfitting. Unlike the pretraining prompts used by MusicGen and Mustango(shown in Table 1)—which are largely Western-centric and lack fine-grained metadata—our prompts explicitly specify genre, melodic structure, instrumentation, and rhythm tailored to Hindustani Classical and Turkish Makam.
Setting Example Prompt
Pretraining-MusicGen “80s pop track with bassy drums and synth”
Pretraining-Mustango “Western instrumental with electric guitar, rim shots, and E major key”
Ours “Hindustani Classical with Tanpura, Tabla, Voice, Harmonium in Bhairavi raga in Teental.”
Table 1. Prompt styles across MusicGen, Mustango, and our approach.
@@ 3.3.2. Subjective Evaluation
For subjective human evaluation, we used a similar framework as described in  Mehta et al. (2025) by implementing an arena-style setup and an elimination-based strategy to compare the generative performance of our models subjectively. In this approach, for each of the 10 distinct prompts, we generated audio snippets from two models at a time and compared them on the basis of clarity, auditory pleasantness, and freedom from unwanted artifacts. The model that wins on a majority of these prompt-wise comparisons advances to the next round, and this process is repeated until a single model remains. We compare five models (three from MusicGen and two from Mustango) for both Hindustani and Makam, resulting in 40 total matchups per genre. Ultimately, the best-performing model is identified as the winner for that genre. For subjective evaluations, the annotators included two people–one music expert and another avid listener who annotated both genres. Due to resource constraints, we limited the number of annotators. Given that our evaluation focused on overall quality rather than a detailed analysis of specific components like melodic line, rhythm, and timbre, we found that two annotators were sufficient.
@@ 4.1. Objective Evaluations
Figure 2 (left plots) shows results for MusicGen for both Hindustani Classical and Turkish Makam music. Across all configurations, we find that the 40M parameter scale offers the best FAD and FD scores ensuring optimal trade-off between music generation quality and the adapter size for the given amount of data. At this scale, FAD and FD scores stabilize at low values across architectures and genres, while training time remains significantly lower than that of the larger 70M models. For example, Hindustani_MG-C and Makam_MG-T at 40M both achieve excellent FAD scores (10.0 and 13.0, respectively), indicating good audio quality, without incurring the diminishing returns observed beyond this scale. This sweet spot can be attributed to the nature of adapter-based PEFT; small adapters (e.g., 2M or 8M) lack sufficient capacity to capture complex musical dependencies, conversely, very large adapters (e.g., 70M) not only increase memory and compute requirements but also risk overfitting or disrupting the model’s learned representations particularly noticeable in MG-L configurations, where FD scores for both genres degrade sharply at 70M. Among architectures, convolution-based adapters (MG-C) performs best for Hindustani classical, while transformer-based adapters (MG-T) perform best for Turkish Makam.
Figure 2 (right plots) shows Mustango results for both Hindustani Classical and Turkish Makam music. Unlike MusicGen, whose best trade-off emerged at the 40M scale Mustango demonstrates its most stable performance in the 20M–40M parameter range. At smaller adapter sizes (2M or 8M), model capacity is insufficient for capturing complex musical structures, reflected in higher FAD scores (e.g., Hindustani_MS-C at 2M hovers near 12). In contrast, the largest 70M adapters yield inconsistent improvements and can even degrade fidelity metrics (e.g., FD for Makam_MS-T climbs above 100 at 70M), suggesting a risk of overfitting or destabilizing previously learned representations. In particular, 40M adapters typically achieve low FAD values, especially for Hindustani CNN (MS-C) at 6.4 and Makam CNN (MS-C) at 8.39. Among architectures for Mustango, convolution-based adapters (MS-C) performs best for both Hindustani classical and Turkish Makam with transformer based adapter (MS-T) matching the performance for Hindustani Classical but not for Turkish Makam.
Both MusicGen and Mustango share the common trend that mid-range adapter sizes are optimal for the trade-offs of too-little capacity (leading to higher FAD/FD scores) and too-large capacity (increased risk of overfitting and larger sizes leading to heavier compute demands). In MusicGen, the optimal point holds firmly at 40M; in Mustango, peak performance is often achieved slightly earlier, around 20M parameters, though 40M remains competitive. In both models, architecture preferences are genre-dependent, with CNN adapters often favored for Hindustani classical and Transformer adapters for Turkish Makam.